OpenStudy (haileyd):

Two balloons, balloon A and balloon B, have the total volume of 3/5 gallons. Balloon A has a greater volume than balloon B. The difference in the volumes is 1/5 gallons. What would be the volume of each balloon?

OpenStudy (haileyd):

@EmilyCheyenne

OpenStudy (haileyd):

@wolfe8

OpenStudy (wolfe8):

A+B=3/5 A=B+1/5 Can you solve that?

OpenStudy (haileyd):

Yah i just didnt know how to put it. Thanks